# Maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$

### v1

@@ 3.1. Maximal discs and Gauss maps
Some remarks are in order about Proposition 3.4. For $(1)\to(2)$, the orientation preserving minimal diffeomorphism $G:\mathbb{D}\to\mathbb{D}$ is obtained by putting $G=G_{+}\circ G_{-}^{-1}$ for the Gauss map $(G_{+},G_{-}):\Sigma\to\mathbb{D}\times\mathbb{D}$. For $(2)\to(4)$, the quasisymmetric homeomorphism $h:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is obtained by restricting $G$ to the boundary $\partial{\mathbb{D}}$. For $(4)\to(1)$, the maximal disc $\Sigma$ is obtained such that $\partial{\Sigma}$ is the graph of $h:\mathbb{S}^{1}\to\mathbb{S}^{1}$. For the equivalence of $(1),(3),(4)$, one can use a similar construction by (3.10) as above instead of $(2)$.
@@
By Proposition 3.4, the followings equivalently represent the same points in the universal Teichmüller space $T(1)$ respectively.
  1. (1)
    the conformal structure of $\mathbb{D}_{\pm}:=(\mathbb{D},z_{\pm})$,
  2. (2)
    the quasisymmetric homeomorphism $z_{\pm}|_{\mathbb{S}^{1}}=(F_{\pm}\circ z)|_{\mathbb{S}^{1}}$,
  3. (3)
    the Beltrami differential $\mu_{z_{\pm}}$ of the map $z_{\pm}$.
For a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, the induced conformal structure on $\Sigma$ determines $(\mathbb{D},z)$, hence the Beltrami differential $\mu$ of $z$. The Hopf differentials of the induced Gauss maps $F_{\pm}$ are ±$\pm\Phi$ by the construction. Hence, the maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ determines a pair $(\mu,\Phi)$ in $T^{*}T(1)$. Now, we introduce the following definition,
@@
For a maximal conformal embedding $\sigma$ of $(\mathbb{D},z)$, the maximal disc ${\Sigma}\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is of Weil-Petersson class if the associate pair $(\mu,\Phi)\in T^{*}T_{0}(1)$.
@@
For a maximal disc of Weil-Petersson class $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ with the induced metric $e^{\phi}|dz|^{2}$ and the induced Gauss map $(F_{+},F_{-}):\Sigma\to\mathbb{D}_{+}\times\mathbb{D}_{-}$, we have
@@
    $\int_{\Sigma}\kappa^{2}e^{\phi}\ d^{2}z=\int_{\Sigma}|\mu_{G_{\pm}}|^{2}e^{\phi}\,d^{2}z=\int_{\Sigma}|\mu_{F_{\pm}}|^{2}e^{\phi}\,d^{2}z<\infty,$
@@
    $\int_{\Sigma}|\Phi(G_{\pm})|^{2}e^{-\phi}\,d^{2}z=\int_{\Sigma}|\Phi(F_{\pm})|^{2}e^{-\phi}\,d^{2}z<\infty.$
@@
By the condition that $\Sigma$ is Weil-Petersson class and Lemma 3.3 of [15],
$\int_{\Sigma}|\mu_{F_{\pm}}|^{2}e^{\psi}\,d^{2}z<\infty.$
This implies that $|\mu_{F_{\pm}}|(z)\to 0$ as $|z|\to 1$. Now, recall that there exists a nonzero constant $a$ such that the curvature $-1<K_{\phi}<-a^{2}$ by (3.6) and $|\mu_{\pm}|\to 0$. Then, we have that $e^{\phi}<a^{-2}e^{\psi}$ by Lemma 4.9 of [21]. By these facts and $\kappa^{2}=|\mu_{F_{\pm}}|^{2}$,
(3.11) $\int_{\Sigma}\kappa^{2}e^{\phi}\ d^{2}z=\int_{\Sigma}|\mu_{F_{\pm}}|^{2}e^{\phi}\,d^{2}z<a^{-2}\int_{\Sigma}|\mu_{F_{\pm}}|^{2}e^{\psi}\,d^{2}z<\infty$
This completes the proof of the item (1). Then the item (2) follows from this easily noticing that $|\Phi(F_{\pm})|^{2}e^{-\phi}=|\mu_{F_{\pm}}|^{2}e^{\phi}$.
@@ 3.2. Mess map
For a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, its first and second fundamental forms have the local expressions in terms of the induced conformal structure on $\Sigma$ and the Hopf differential $\Phi(G_{\pm})$ as in (3.2) and (3.4). Then, each component of the Gauss map $G_{\pm}$ determines the conformal structures $(\mathbb{D},z_{\pm})$ respectively. Denoting by $I_{\pm}$ the hyperbolic metric on $(\mathbb{D},z_{\pm})$, we have
(3.12) $I_{\pm}=I(E\pm JI^{-1}II\cdot\,,E\pm JI^{-1}II\cdot\,)$
where $E$ denotes the identity map and $J$ denotes the complex structure on $T\Sigma$. The equality (3.12) is proved in Lemma 3.16 of [8]. Note that the equality (3.12) makes sense originally for the pair $(I,II)$ given by $(\mu,\Phi)$ as in (3.2) and (3.4), which is associated to a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. In particular, the Beltrami differential $\mu$ corresponds to the induced conformal structure on $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$.
Naturally, we can extend the above correspondence given in (3.12) for any $(\mu,\Phi)\in T^{*}T(1)$ where $\mu$ may not correspond to the induced conformal structure on $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. For a $(\mu,\Phi)\in T^{*}T(1)$, we solve the Gauss equation (3.5) with respect to $\Phi$ on $(\mathbb{D},z)$ where $z$ denotes the coordinate corresponding to $\mu$. Then we define by $I(\mu,\Phi):=e^{\phi}|dz|^{2}$ where $\phi$ denotes the solution of the Gauss equation. Moreover, by Theorem 3.2 of [18], there exist harmonic maps
$F_{\pm}:(\mathbb{D},z)\to(\mathbb{D},z_{\pm}),$
which we also call induced Gauss map in this general case. Now, we introduce the Mess map defined by
@@
which maps $(\mu,\Phi)$ to $(I_{+},I_{-})$ using (3.12) for $I(\mu,\Phi)$ and $II$ given by (3.4). By the result in the section 5 of [7], the map $\mathrm{Mess}$ is bijective.
Remark 3.8.
By the above constructions of the induced Gauss map $F_{\pm}$, one can see that the one-to-one correspondence of the items (3) and (4) in Proposition 3.4 still holds in this general situation.
Remark 3.9.
For a totally geodesic maximal disc $(\mathbb{D},w)\cong\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, we can see that $I_{\pm}=I=e^{\psi(w)}|dw|^{2}$ by (3.12) since $II=0$. On the other hand, for a pair $(\mu,0)\in T^{*}T(1)$ representing a different conformal structure $(\mathbb{D},z)$ on a (real) totally geodesic maximal disc $\Sigma$, the map $\mathrm{Mess}$ sends $(\mu,0)\in T^{*}T(1)$ to $I_{\pm}=e^{\psi(z)}|dz|^{2}$, which is not the induced metric $e^{\psi(w)}|dw|^{2}$. Hence, the image of the induced Gauss map $(F_{+},F_{-})$ of the totally geodesic disc $(\mathbb{D},z)$ is always realized as the diagonal minimal Lagrangian submanifold in $(\mathbb{D},z_{+})\times(\mathbb{D},z_{-})$ where both discs are equipped with the hyperbolic metrics. This also implies that the map $\mathrm{Mess}$ sends the image of zero section of $T^{*}T(1)$ to the diagonal set in $T(1)\times T(1)$.
Restricting the Mess map to the holomorphic cotangent bundle of the Weil-Petersson Teichmüller space $T_{0}(1)$, we have
@@
The following restriction of $\mathrm{Mess}$ is a diffeomorphism,
@@
Recall that a point in $T^{*}T_{0}(1)$ is represented by a pair $(\mu,\Phi)$ for a Beltrami differential $\mu$ on $\mathbb{D}$ and a holomorphic quadratic differential $\Phi$ on $\mathbb{D}$ satisfying the conditions

### v2

@@ 3.1. Maximal discs and Gauss maps
Here are some remarks regarding Proposition 3.4:
  • (1) $\to$ (2): The orientation-preserving minimal Lagrangian diffeomorphism $G:\mathbb{D}_{w}\to\mathbb{D}_{w}$ is obtained by defining
    $G=G_{+}\circ G_{-}^{-1}$
    where $(G_{+},G_{-}):\mathbb{D}_{z}\to\mathbb{D}_{w}\times\mathbb{D}_{w}$ is the Gauss map of the maximal disc $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ (see Theorem 4.1 of [2] for more details).
  • (2) $\to$ (3): The harmonic maps $F_{\pm}$ can be obtained from $G_{\pm}$ as previously explained (see Lemma 2.1 of [2]).
  • (3) $\to$ (4): The quasisymmetric homeomorphism $h:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is obtained by restricting $F$ to the boundary $\partial\mathbb{D}$, i.e.,
    $h:=F|_{\partial\mathbb{D}}=G|_{\partial\mathbb{D}}.$
    Here the identification of $\partial{\mathbb{D}_{z_{-}}}$ with $\partial{\mathbb{D}}$ is made via the restriction of the coordinate map of $\mathbb{D}_{z_{-}}$.
  • (4) $\to$ (1): Given a quasisymmetric homeomorphism $h:\mathbb{S}^{1}\to\mathbb{S}^{1}$, the corresponding maximal disc $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is constructed such that its boundary $\partial\Sigma=\partial(\sigma(\mathbb{D}_{z}))$ is the graph of $h$ (see Theorem 1.10 of [5] for more details).
By Remark 5.12 of [6], for an orientation preserving minimal Lagrangian diffeomorphism $G:\mathbb{D}_{z}\to\mathbb{D}_{z}$ and the hyperbolic metric $g_{z}$ determining the conformal structure of $\mathbb{D}_{z}$, there exists a $g_{z}$-self-adjoint endomorphism $b\in\mathrm{End}(T\mathbb{D})$ satisfying the following conditions:
(3.11) $G^{*}g_{z}=g_{z}(b\cdot,b\cdot),\qquad d^{\nabla}b=0,\qquad\mathrm{det}\,b=1$
where $\nabla$ denotes the Levi-Civita connection of $g_{z}$. Now, we obtain
Proposition 3.5.
Let $G:\mathbb{D}_{z}\to\mathbb{D}_{z}$ be an orientation-preserving diffeomorphism satisfying the condition (3.11). Consider the map $\sigma_{G,b}:\mathbb{D}_{z}\to\mathrm{Isom}(\mathbb{D}_{z})$ defined such that for $x\in\mathbb{D}$, $\sigma_{G,b}(z(x))$ is the unique isometry
$\gamma\in\mathrm{Isom}(\mathbb{D}_{z})\cong\mathrm{PSL}(2,\mathbb{R})\cong\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$
satisfying the following conditions:
(3.12) $\gamma\circ G(z(x))=z(x),\qquad d\gamma_{G(z(x))}\circ dG_{z(x)}=-b_{z(x)}.$
Then, the following equality holds:
(3.13) $\sigma_{G,b}^{*}(g_{\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}})=\frac{1}{4}g_{z}((E+b)\cdot,(E+b)\cdot)$
where $E$ denotes the identity on $\mathrm{End}(T\mathbb{D})$.
Proof.
The claim for the case of $\mathbb{D}_{w}$ is established in Proposition 5.5 of [6]. Now, for a given diffeomorphism $G:\mathbb{D}_{z}\to\mathbb{D}_{z}$, consider the following commutative diagram:
(3.14) ${\mathbb{D}_{w}}$${\mathbb{D}_{w}}$${\mathbb{D}_{z}}$${\mathbb{D}_{z}}$$\scriptstyle{z}$$\scriptstyle{\widetilde{G}}$$\scriptstyle{z}$$\scriptstyle{G}$
where $\widetilde{G}$ and $G$ are diffeomorphisms over $\mathbb{D}_{w}$ and $\mathbb{D}_{z}$ respectively, satisfying the given conditions in (3.11). By assumption, the diffeomorphism ${G}:\mathbb{D}_{z}\to\mathbb{D}_{z}$ also satisfies the conditions in (3.12). In particular, for $x\in\mathbb{D}$ , the map ${\sigma}_{{G},{b}}$ maps $(z(x))$ to an isometry ${\gamma}\in\mathrm{Isom}(\mathbb{D}_{z})\cong\mathrm{PSL}(2,\mathbb{R})$ that satisfies the equalities in (3.12). From (3.14), note that the identification between $\mathrm{Isom}(\mathbb{D}_{z})$ and $\mathrm{Isom}(\mathbb{D}_{w})$ is given by
$\gamma\in\mathrm{Isom}(\mathbb{D}_{z})\to\tilde{\gamma}:=z^{-1}\circ{\gamma}\circ z\in\mathrm{Isom}(\mathbb{D}_{w}).$
Since ${\sigma}_{{G},{b}}(z(x))\in\mathrm{Isom}(\mathbb{D}_{z})\cong\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ for $x\in\mathbb{D}_{w}$, it follows that
$\tilde{\gamma}:=\tilde{\sigma}_{\widetilde{G},\tilde{b}}(x)=z^{-1}\circ{\sigma}_{{G},{b}}(z(x))\circ z=z^{-1}\circ{\gamma}\circ z.$
We define $\tilde{b}\in\mathrm{End}(T\mathbb{D})$ by
(3.15) $\tilde{b}=dz^{-1}\circ{b}\circ dz,$
which satisfies (3.11) and (3.12) for Levi-Civita connection of $g_{w}$. Here, the equality $d^{\widetilde{\nabla}}\tilde{b}=0$ follows from the corresponding equality for ${b}$ and the relation
~$\widetilde{\nabla}=dz^{-1}\circ{\nabla}\circ dz^{-1},$
[…]
Applying Proposition 5.5 of [6] under the above conditions for the special case $\mathbb{D}_{w}$, the following equality holds
(3.16) $\tilde{\sigma}_{\widetilde{G},\tilde{b}}^{*}(g_{\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}})=\frac{1}{4}g_{w}((E+\tilde{b})\cdot,(E+\tilde{b})\cdot).$
Then, finally, we obtain
$\sigma_{{G},{b}}^{*}(g_{\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}})=(z^{-1})^{*}(\tilde{\sigma}_{\widetilde{G},\tilde{b}})^{*}(g_{\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}})=\frac{1}{4}(z^{-1})^{*}\big{(}g_{w}((E+\tilde{b})\cdot,(E+\tilde{b})\cdot)\big{)}\\ =\frac{1}{4}g_{w}((E+\tilde{b})dz^{-1}\cdot,(E+\tilde{b})dz^{-1}\cdot)=\frac{1}{4}g_{w}(dz^{-1}(E+dz\,\tilde{b}\,dz^{-1})\cdot,dz^{-1}(E+dz\,\tilde{b}\,dz^{-1})\cdot)\\ =\frac{1}{4}(z^{-1})^{*}g_{w}((E+dz\,\tilde{b}\,dz^{-1})\cdot,(E+dz\,\tilde{b}\,dz^{-1})\cdot)=\frac{1}{4}(g_{z}((E+{b})\cdot,(E+{b})\cdot).$
This completes the proof. ∎
Hence, by Propositions 3.4 and 3.5, the first fundamental form $I$ over the maximal conformal embedding ${\sigma}:\mathbb{D}_{w}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is given by
(3.17) $I=\frac{1}{4}g_{w}((E+b)\cdot,(E+b)\cdot),$
where $b\in\mathrm{End}(T\mathbb{D})$ satisfies conditions (3.11) and (3.12) for $g_{w}$.
@@
We apply Proposition 3.5 in the special case where $b=0$. This condition implies that $G$ is the identity map on $\mathbb{D}_{z}$. Hence, the equalities in (3.12) indicates that $\gamma\in\mathrm{PSL}(2,\mathbb{R})$ represents an involutional rotation by an angle $\pi$ around the point $z(x)$, which we denote by $\mathcal{I}_{z(x)}$. Therefore, the image of $\sigma_{G,b}:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is the totally geodesic disc
(3.18) $\mathcal{R}_{\pi}:=\{\,\mathcal{I}_{z(x)}:x\in\mathbb{D}\,\}\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1},$
which can be identified with $\mathbb{D}_{w}$. In this setting, the map $\sigma_{G,b}:\mathbb{D}_{z}\to\mathcal{R}_{\pi}\cong\mathbb{D}_{w}$ coincides the map $w_{\mu}$ where $\mu$ is the Beltrami differential of $z$. Recall that $w_{\mu}:\mathbb{D}_{z}\to\mathbb{D}_{w}$ is a conformal map. Additionally, in this case, we observe that $G_{+}=G_{-}$ is the identity map on $\mathbb{D}_{w}$, leading to the identification $\mathbb{D}_{z_{+}}=\mathbb{D}_{z_{-}}=\mathbb{D}_{z}$.
@@
Let ${\sigma}:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ be a maximal conformal embedding such that the boundary $\partial\Sigma=\partial(\sigma(\mathbb{D}_{z}))$ is the graph of a quasisymmetric homeomorphism $h$ representing an element in $T(1)$. We say that the maximal conformal embedding ${\sigma}:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is of Weil-Petersson class if the corresponding quasisymmetric homeomorphism $h$ represents an element in $T_{0}(1)$.
@@
For a maximal conformal embedding of Weil-Petersson class
$\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$
with the induced metric $e^{\phi}|dz|^{2}$ and the induced Gauss map $(F_{+},F_{-}):\mathbb{D}_{z}\to\mathbb{D}_{z_{+}}\times\mathbb{D}_{z_{-}}$, we have:
@@
    $\int_{\mathbb{D}}\kappa^{2}e^{\phi}\ d^{2}z=\int_{\mathbb{D}}|\mu_{G_{\pm}}|^{2}e^{\phi}\,d^{2}z=\int_{\mathbb{D}}|\mu_{F_{\pm}}|^{2}e^{\phi}\,d^{2}z<\infty.$
@@
    $\int_{\mathbb{D}}|\Phi(G_{\pm})|^{2}e^{-\phi}\,d^{2}z=\int_{\mathbb{D}}|\Phi(F_{\pm})|^{2}e^{-\phi}\,d^{2}z<\infty.$
@@
Note that the quasisymmetric homeomorphism $h:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is given by the restriction of $F:=F_{+}\circ F_{-}^{-1}$ to the boundary $\partial\mathbb{D}$. Then, by the condition that $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is of Weil-Petersson class and using Lemma 3.3 of [18], we have the following:
(3.19) $\int_{\mathbb{D}}|\mu_{F}|^{2}e^{\psi}\,d^{2}z_{-}<\infty.$
Now, form (2.10), the Beltrami differential $\mu_{F}$ can be expressed as:
where $\mu_{\pm}=\mu_{F_{\pm}}$. This leads to the expression:
Furthermore, we have
(3.22) $\int_{\mathbb{D}}|\mu_{F}|^{2}e^{\psi}\,d^{2}z_{-}=\,\int_{\mathbb{D}}\left(|\mu_{F}|^{2}\circ F_{-}\right)\,e^{\psi\circ F_{-}}\,\left(\left|(F_{-})_{z}\right|^{2}-\left|(F_{-})_{\bar{z}}\right|^{2}\right)\,d^{2}z\\ =\,\int_{\mathbb{D}}\left(|\mu_{F}|^{2}\circ F_{-}\right)\,e^{\psi\circ F_{-}}\,\left|(F_{-})_{z}\right|^{2}\left(1-\left|\mu_{-}\right|^{2}\right)\,d^{2}z\\ =\,\int_{\mathbb{D}}\frac{4|\mu_{-}|^{2}}{(1+|\mu_{-}|^{2})^{2}}e^{\phi}\left(1-\left|\mu_{-}\right|^{2}\right)\,d^{2}z.$
Since $[\mu_{F}]\in T_{0}(1)$, we have $|\mu_{F}(z)|\to 0$ as $|z|\to 1$, implying that $|\mu_{-}(z)|\to 0$ as $z$ approaches to $\partial\mathbb{D}$ by (3.21). Hence, there exists $C>0$ such that
(3.23) $C<\frac{1-|\mu_{-}|^{2}}{(1+|\mu_{-}|^{2})^{2}}<1\qquad\text{over}\quad\mathbb{D}.$
Hence, from (3.22) and (3.23), we conclude that
(3.24) $\int_{\mathbb{D}}|\mu_{F}|^{2}e^{\psi}\,d^{2}z_{-}<\infty\qquad\text{ if and only if}\qquad\int_{\mathbb{D}}{|\mu_{\pm}|^{2}}e^{\psi}\,d^{2}z\leq\int_{\mathbb{D}}{|\mu_{\pm}|^{2}}e^{\phi}\,d^{2}z<\infty.$
This means that $F_{\pm}|_{\partial{\mathbb{D}}}$ represents points in $T_{0}(1)$, which also implies that $|\mu_{F_{\pm}}|(z)\to 0$ as $|z|\to 1$. Now, recall that there exists a nonzero constant $a$ such that the curvature satisfies
$-1<K_{\phi}<-a^{2}$
by (3.6) and $|\mu_{\pm}|\to 0$. By Lemma 4.9 of [24], it follows that
$e^{\phi}<a^{-2}e^{\psi}.$
Using these facts and the identity $\kappa^{2}=|\mu_{F_{\pm}}|^{2}$, we obtain
(3.25) $\int_{\mathbb{D}}\kappa^{2}e^{\phi}\ d^{2}z=\int_{\mathbb{D}}|\mu_{F_{\pm}}|^{2}e^{\phi}\,d^{2}z<a^{-2}\int_{\mathbb{D}}|\mu_{F_{\pm}}|^{2}e^{\psi}\,d^{2}z<\infty.$
This completes the proof of item (1). The proof of item (2) follows easily by noting that
$|\Phi(F_{\pm})|^{2}\,e^{-\phi}=|\mu_{F_{\pm}}|^{2}\,e^{\phi}.$
@@ 3.2. Mess map
For a given pair $(\mu,\Phi)$ representing a point in the holomorphic cotangent bundle $T^{*}T(1)$, we first take a conformal structure over the unit disc determined by $\mu$, which we denote by $\mathbb{D}_{z}$. For a given the holomorphic quadratic differential $\Phi$ over $\mathbb{D}_{z}$ and $\mathbb{D}_{w}$, by Theorem 3.2 of [21], there exist harmonic maps
$G_{\pm}:\mathbb{D}_{z}\to\mathbb{D}_{w}$
with the Hopf differential ±$\pm\Phi$ respectively. Then, the pullback by $G_{\pm}$ of the hyperbolic metric on $\mathbb{D}_{w}$ are two different hyperbolic metrics on $\mathbb{D}_{z}$, denoted by $I_{\pm}$, which determines two conformal structures $\mathbb{D}_{z_{\pm}}$ respectively. As explained previously, there exist harmonic maps
$F_{\pm}:\mathbb{D}_{z}\to\mathbb{D}_{z_{\pm}}$
with the Hopf differentials ±$\pm\Phi$ respectively. Finally, we introduce the Mess map defined as
@@
which maps a pair $(\mu,\Phi)$ to the pair $(I_{+},I_{-})$ representing a point in $T(1)\times T(1)$.
Proposition 3.9.
The map $\mathrm{Mess}:T^{*}T(1)\to T(1)\times T(1)$ is a bijective map.
Proof.
To show the injectivity of $\mathrm{Mess}:T^{*}T(1)\to T(1)\times T(1)$, we consider the following equalities
where $\mu_{+}=\mu_{F_{+}}$. If the $\mathrm{Mess}$ map is not injective, there are two pairs $(a,b)$ and $(a^{\prime},b^{\prime})$ for $(\mu_{z},z^{*}(\mu_{F}))$ that satisfy the same equations for a given $\mu_{z_{\pm}}$, leading to the equalities:
By rearranging and combing these expressions, we obtain
(3.29) $a-a^{\prime}=(\bar{a}-\bar{a^{\prime}})bb^{\prime}.$
From this, if $a\neq a^{\prime}$, we have $|bb^{\prime}|=1$. However, this contradicts the fact that $|bb^{\prime}|<1$. Hence, we conclude that $a=a^{\prime}$, which further implies that $b=b^{\prime}$ by the previous equation (3.28). Thus, there exists a unique pair $(\mu_{z},\mu_{F_{+}})$ satisfying (3.27) for a given pair $(\mu_{z_{+}},\mu_{z_{-}})$. Since the pair $(\mu_{z},\mu_{F_{+}})$ uniquely determines the pair $(\mu_{z},\Phi)$ by (3.9), this proves the injectivity of the map $\mathrm{Mess}:T^{*}T(1)\to T(1)\times T(1)$.
To prove the surjectivity of the map $\mathrm{Mess}:T^{*}T(1)\to T(1)\times T(1)$, let $(I_{+},I_{-})$ be a pair representing a point in $T(1)\times T(1)$ be a given pair. We first consider two conformal structures over the unit disc determined by $I_{+},I_{-}$, denote by $\mathbb{D}_{z_{+}}$ and $\mathbb{D}_{z_{-}}$, respectively. This implies that there exist two quasiconformal maps $z_{\pm}:\mathbb{D}_{w}\to\mathbb{D}_{z_{\pm}}$. Next, we consider the conformal structure on the unit disc $\mathbb{D}$ determined by $I_{+}+I_{-}$, which we denote by $\mathbb{D}_{z}$. By applying Proposition 3.4, for the fixed $\mathbb{D}_{z}$ and the quasisymmetric homeomorphism $(z_{+}\circ z_{-}^{-1})|_{\partial\mathbb{D}}$, there exists a maximal conformal embedding $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ with its induced Gauss map $(F_{+},F_{-}):\mathbb{D}_{z}\to\mathbb{D}_{z_{+}}\times\mathbb{D}_{z_{-}}$. By the construction, the pair $(\mu_{z},\Phi(F_{+}))$ is mapped to the given pair $(I_{+},I_{-})$ by the map $\mathrm{Mess}$. This completes the proof of the surjectivity of $\mathrm{Mess}$.
When restricting the Mess map to the holomorphic cotangent bundle of the Weil-Petersson Teichmüller space $T_{0}(1)$, we have
@@
The following restriction of the $\mathrm{Mess}$ map is a diffeomorphism,
@@
Recall that a point in $T^{*}T_{0}(1)$ is represented by a pair $(\mu,\Phi)$, where $\mu$ is a Beltrami differential on $\mathbb{D}_{w}$ and $\Phi$ is a holomorphic quadratic differential on $\mathbb{D}_{z}$ satisfying the conditions